A particle executes SHM in accordance with x=Asin (wt). If t1 is the time taken by it to reach from x=0 to x=(A/2)√3 and t2 is the time taken by it to reach from x=(A/2)√3 to x=A, the value of t1/t2 is

Grade:12

1 Answers

Khimraj
6 years ago
x = Asin(wt)
t1 will be given by 
(A/2)√3 = Asin(wt1)
So sin(wt1) = √3/2
wt1 = π/3   – (1)
time taken by particle to reach from x=0 to x=A will be t1 +t2.
So
A = A sinw(t1+t2)
So wt1 + wt2 = π/2
then wt2 = π/2 – π/3
So wt2 = π/6  – (2)
divide equation (1) by (2)
t1/t2 = 2.
Hope it clears.
If you like anser then please approve it.